# Adaptive Deadline and Batch Layered Synchronized Federated Learning

### v1

@@ 3.1 Optimization Problem
Section (2.3) motivates formulating a constrained optimization problem holding R1-R3. To cope with R3, we allow the batch sizes to vary between users, drawing inspiration from previous works [18, 19] that advocated the usage of different batch sizes to deal with heterogeneity. Accordingly, we commence by determining a cost function that captures the model performance as a function of the deadlines $\{T_{t}^{d}\}_{t=1}^{R}$ and the local batch sizes $\{S_{u}\}_{u=1}^{U}$. To maintain a tractable analysis, we use a single parameter $m$ to control the batch size across devices. Specifically, we introduce a parameter $P_{u}$ for each user $u$, that is related to its computational capability, and set its batch size to be $S_{u}=\left\lceil m\cdot P_{u}\right\rceil$. This proportionality ensures that faster devices do not dominate training by contributing excessively more updates within the same deadline, which could bias the model toward their data and hinder convergence to the global optimum ${\bm{w}}_{\rm opt}$ in (1).
@@
Following the common practice in the FL literature [6, 20, 21], we assess the learned FL model via its expected $\ell_{2}$ distance to the optimal FL model. Accordingly, by letting $\tilde{{\bm{w}}}_{t}\left(\{T_{\tau}^{d}\},m\right)$ denote the global model obtained at round $t$ via layer-wise aggregation (5) with deadlines $\{T_{\tau}^{d}\}_{\tau\leq t}$ and mini-batch sizes $\{S_{u}=\left\lceil m\cdot P_{u}\right\rceil\}_{u=1}^{U}$, we design ADEL-FL based on the following optimization framework:
@@
$\min_{\{T_{t}^{d}\}_{t=1}^{R},m}$ $\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}\left(\{T_{t}^{d}\}_{t=1}^{R},m\right)-{\bm{w}}_{\rm opt}\right\|^{2}\right]\quad\textrm{s.t.}\ \sum_{t=1}^{R}T_{t}^{d}\leq T_{\max}.$
3.2 Deriving the Cost Function
@@
3.2.1 Assumptions
We perform the derivation of the cost function under the following assumptions, commonly adopted in FL studies [6, 22, 11, 23]:
@@ 3.2.1 Assumptions
    $\begin{gathered}\left({\bm{w}}_{1}-{\bm{w}}_{2}\right)^{T}\nabla F_{u}\left({\bm{w}}_{2}\right)+\frac{1}{2}\rho_{c}\left\|{\bm{w}}_{1}-{\bm{w}}_{2}\right\|^{2}\leq F_{u}\left({\bm{w}}_{1}\right)-F_{u}\left({\bm{w}}_{2}\right)\leq\\ \left({\bm{w}}_{1}-{\bm{w}}_{2}\right)^{T}\nabla F_{u}\left({\bm{w}}_{2}\right)+\frac{1}{2}\rho_{s}\left\|{\bm{w}}_{1}-{\bm{w}}_{2}\right\|^{2}.\end{gathered}$
@@
    Variance Bound: For each user $u$ and round index $t$, the variance of the batch gradient $\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)$ is bounded by $\frac{\sigma_{u}^{2}}{S_{u}}$ for all ${\bm{w}}\in\mathbb{R}^{n}$, i.e.,
    $\mathbb{E}\left[\left\|\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)-\nabla F_{u}({\bm{w}})\right\|^{2}\right]\leq\frac{\sigma_{u}^{2}}{S_{u}}.$
@@
    Gradient Bound: For each user $u$ and round index $t$, the expected squared $\ell_{2}$ norm of the stochastic gradients $\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)$ is uniformly bounded by some $G^{2}$ for all ${\bm{w}}\in\mathbb{R}^{m}$, i.e.,
@@
  4. A4
    Computational Model: Let $T_{t,u}^{b,l}$ be the computation time of user $u$ in round $t$ for the backpropagation of layer $l$. The random variables $\{T_{t,u}^{b,l}\}$ are i.i.d (in $u$, $t$, and $l$) and obey an exponential distribution:
    $T_{t,u}^{b,l}\sim{\rm Exp}\left(\frac{S_{u}}{P_{u}}\right).$
Under assumption A2, the variance of the batch gradient is inversely proportional to the batch size, referring to sampling with replacement of the samples. This relation is crucial in FL scenarios, as it directly links the user-level batch size to stability and convergence guarantees. Such variance-batch proportionality is supported in prior works (e.g., [15], [6]). Assumption A4 models the computation time of a layer as an exponential random variable. The exponential model is memoryless and well-suited for modeling compute time variability in deadline-constrained execution. This assumption aligns with prior work in straggler modeling and distributed training analysis [24, 25, 26], where exponential or shifted-exponential distributions are used to enable tractable performance bounds and scheduling policies. In addition, we define the heterogeneity gap, quantifying the differences in data distribution between users [6]:
$\Gamma\triangleq F\left({\bm{w}}_{\mathrm{opt}}\right)-\frac{1}{U}\sum_{u=1}^{U}\min_{{\bm{w}}}F_{u}({\bm{w}}),$
where ${\bm{w}}_{\rm opt}$ is defined in (1).
3.2.2 Convergence Bound
After presenting the assumptions, we can now derive the bias correction factor $p_{l}$ defined in (5):
@@ 3.2.2 Convergence Bound
When Assumptions A1-A4 hold, the value of $p_{l}\triangleq P(|\mathcal{U}_{t}^{l}|=0)$ is given by:
$p_{l}=Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U},$
[…]
As expected, $p_{l}$ is monotonically decreasing with the layer index $l$, as the recursive structure of the backpropagation algorithm implies that first layers are less likely to be reached under tight deadline. Using Lemma 1, we state two lemmas for unbiasedness and bounded variance of each round of our proposed framework:
@@
(Unbiasedness). Under Assumptions A1-A4, for every FL round of index $t$, the global model aggregated via ADEL-FL using (5) is an unbiased estimator of the one obtained via vanilla FedAvg [6], namely,
$\mathbb{E}\left[\tilde{{\bm{w}}}_{t+1}^{l}\right]={\bm{w}}_{t+1}^{l},$
where ${\bm{w}}_{t+1}^{l}$ is given by (3).
@@
(Bounded variance) Consider ADEL-FL with Assumptions A1-A4 and the learning rate $\eta_{t}$ set to be non-increasing and satisfying $\eta_{t}\leq 2\eta_{t+1}$ and $p_{1}<0.5$. Then, for every FL round $t$, the expected difference between $\tilde{{\bm{w}}}_{t+1}$ and ${\bm{w}}_{t+1}$, which is the variance of $\tilde{{\bm{w}}}_{t+1}$, is bounded by:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]\leq\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}{1-2Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}.$
Lemmas 2 and 3 are critical for establishing the convergence bound. Lemma 2 ensures that ADEL-FL update remains an unbiased estimator of the FedAvg update, preserving convergence direction. Lemma 3 guarantees that the variance introduced by partial updates remains controlled. These properties mirror the core assumptions required in classical FL convergence analyses, such as those in [6], and are necessary to derive a meaningful upper bound on the expected distance to ${\bm{w}}_{\rm opt}$.
We are now ready to state the main theoretical result of this work: the cost function of Problem 1, which is the convergence bound for the ADEL-FL algorithm under Assumptions A1-A4. This result characterizes how the joint selection of training deadlines and batch size scaling affects the convergence behavior, accounting for both user heterogeneity and partial gradient contributions.
@@
Consider ADEL-FL with Assumptions A1-A4, where the learning rate $\eta_{t}$ is set to be non-increasing and holds $\eta_{t}\leq 2\eta_{t+1}$ and $\eta_{1}\leq\frac{1}{4\rho_{s}}$. Then, we have:
$\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}-{\bm{w}}_{\rm opt}\right\|^{2}\right]\leq\prod_{t=1}^{R}\left(1-\eta_{t}\rho_{c}\right)\Delta_{1}+\sum_{t=1}^{R}\eta_{t}^{2}\left(B+C_{t}\right)\prod_{\tau=t+1}^{R}\left(1-\eta_{\tau}\rho_{c}\right),$
where $\Delta_{1}\triangleq\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{1}-{\bm{w}}_{opt}\right\|^{2}\right]$, and $B$ and $C_{t}$ are given by:
$B\triangleq\frac{1}{mU^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{P_{u}}+6\rho_{s}\Gamma\quad C_{t}\triangleq G^{2}\frac{4U}{U-1}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}{1-2Q\left(L+1-l,\frac{T_{t}^{d}}{m}\right)^{U}}.$
The convergence bound depends on the constants $B$ and $C_{t}$, which capture a trade-off between SGD variance reduction and deadline-induced truncation. This trade-off is analyzed in detail in Section 3.4.
3.3 ADEL-FL Algorithm
Having derived the cost function in Theorem 1, we now formulate the full optimization problem, which jointly determines the round-wise deadlines $\{T_{t}^{d}\}_{t=1}^{R}$ and the global batch-scaling parameter $m$ based on Problem 1. This optimization reflects the practical trade-off between computation progress and user participation in time-limited heterogeneous federated environments.
@@ 3.3 ADEL-FL Algorithm
We solve Problem 2 using a trust-region method [28], an iterative strategy well-suited for nonconvex constrained optimization. Trust-region approaches approximate the objective by a local quadratic model and restrict updates to a neighborhood where the model is deemed reliable. The region’s size is adaptively adjusted based on how well the approximation predicts true improvement. The optimization yields both the per-round deadlines and the batch-scaling factor. These are then embedded into the ADEL-FL procedure, summarized as Algorithm 1. Initially (Steps 1-3), the server solves the global optimization problem to allocate per-round deadlines $T_{t}^{d}$ and computes each user’s batch size $S_{u}=\lceil m\cdot P_{u}\rceil$ proportional to their computational capability. In each round $t$, users perform local training within their time limit (Steps 5-7), computing gradients up to a layer $d_{t}^{u}$ based on how far they progress before the deadline, and then transmit those partial updates to the server. The server aggregates received gradients in a layer-wise manner (Steps 8-11), forming the global model update via the rule (5) while preserving the last available state when no updates arrive for a layer. This procedure repeats for $R$ rounds, and the final model is returned in Step 14.
@@
1:Input: Number of rounds $R$, training deadline $T_{\max}$, learning rates $\left\{\eta_{t}\right\}_{t=1}^{R}$, computational capabilities $\{P_{u}\}_{u=1}^{U}$ and initial model ${\bm{w}}_{0}$
2:    Solve Problem 2 to find $\{T_{t}^{d}\}_{t=1}^{R},m$
3:    Define batch sizes $S_{u}=\left\lceil m\cdot P_{u}\right\rceil$ for each $u\in\{1,\ldots,U\}$
@@
5:         Clients Side - do in parallel for each $u$, until deadline $T_{t}^{d}$:
@@
8:         Server Side - do after deadline time $T_{t}^{d}$:
@@
The overall strategy of ADEL-FL refines the convergence guarantees derived earlier by optimally tuning the time-budgeted training configuration. By solving Problem 2, we adaptively allocate deadlines and batch sizes to each round, enabling balanced participation and efficient learning. The resulting algorithm demonstrates improved performance under strict time constraints, as empirically demonstrated in Section 4.
3.4 Discussion

### v2

@@ III-A Optimization Problem
The formulation in Section II-C motivates formulating a constrained optimization problem holding R1-R3. To cope with R3, we allow the batch sizes to vary between users, drawing inspiration from previous works [37, 38] that advocated the usage of different batch sizes to deal with heterogeneity. Accordingly, we commence by determining a cost function that captures the model performance as a function of the deadlines $\{T_{t}^{\rm d}\}_{t=1}^{R}$ and the local batch sizes $\{S_{t}^{u}\}_{u=1,t=1}^{U,R}$. To maintain tractability, we control all user batch sizes $\{S_{t}^{u}\}_{u=1,t=1}^{U,R}$ via a single global scaling parameter $m$. This parameter relates the local workload of each device to the round deadline, and will be formally defined in Section III-B.
@@
Following the common practice in the FL literature [18, 39, 40], assess the learned FL model via its expected $\ell_{2}$ distance to the optimal FL model as defined in Eq. (1). Accordingly, by letting $\tilde{{\bm{w}}}_{t}\left(\{T_{\tau}^{d}\},m\right)$ denote the global model obtained at round $t$ via layer-wise aggregation (5) with deadlines $\{T_{\tau}^{d}\}_{\tau\leq t}$ and mini-batch sizes $\{S_{t}^{u}\}_{u=1,t=1}^{U,R}$, we design ADEL-FL based on the following optimization framework:
@@
$\min_{\{T_{t}^{\rm d}\}_{t=1}^{R},m}\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}\left(\{T_{t}^{\rm d}\}_{t=1}^{R},m\right)-{\bm{w}}_{\rm opt}\right\|^{2}\right]$
$\textrm{s.t.}\sum_{t=1}^{R}T_{t}^{\rm d}\leq T_{\max}.$
III-B Deriving the Cost Function
@@
III-B1 Assumptions
We perform the derivation of the cost function under the following three assumptions, commonly adopted in FL studies [18, 5, 31]:
@@ III-B1 Assumptions
    $\begin{gathered}\left({\bm{w}}_{1}-{\bm{w}}_{2}\right)^{T}\nabla F_{u}\left({\bm{w}}_{2}\right)+\frac{1}{2}\rho_{c}\left\|{\bm{w}}_{1}-{\bm{w}}_{2}\right\|^{2}\\ \leq F_{u}\left({\bm{w}}_{1}\right)-F_{u}\left({\bm{w}}_{2}\right)\leq\\ \left({\bm{w}}_{1}-{\bm{w}}_{2}\right)^{T}\nabla F_{u}\left({\bm{w}}_{2}\right)+\frac{1}{2}\rho_{s}\left\|{\bm{w}}_{1}-{\bm{w}}_{2}\right\|^{2}.\end{gathered}$
@@
    Variance Bound: For each user $u$ and round index $t$, the variance of the batch gradient $\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)$ is bounded by $\frac{\sigma_{u}^{2}}{S_{t}^{u}}$ for all ${\bm{w}}\in\mathbb{R}^{n}$, i.e.,
    $\mathbb{E}\left[\left\|\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)-\nabla F_{u}({\bm{w}})\right\|^{2}\right]\leq\frac{\sigma_{u}^{2}}{S_{t}^{u}}.$
@@
    Gradient Bound: For each user $u$ and round index $t$, the expected squared $\ell_{2}$ norm of the stochastic gradients $\nabla F_{u}\left({\bm{w}};i_{t}^{u}\right)$ is uniformly bounded by some $G^{2}$ for all ${\bm{w}}\in\mathbb{R}^{n}$, i.e.,
@@
Under Assumption A2, the variance of the batch gradient is inversely proportional to the batch size, referring to sampling with replacement of the samples. This relation is crucial in FL scenarios, as it directly links the user-level batch size to stability and convergence guarantees. Such variance-batch proportionality is supported in prior works (e.g., [5, 18]). Beyond these fundamental analytical assumptions, we also specify the following modeling regarding batch sizes and computational latency, which are grouped separately to distinguish them from the core function properties:
  1. B1
    Computational Model: Let $T_{t,l}^{{\rm b},u}$ be the computation time of user $u$ in round $t$ for the backpropagation of layer $l$. The random variables $\{T_{t,l}^{{\rm b},u}\}$ are i.i.d (in $u$, $t$, and $l$) and obey an exponential distribution:
    $T_{t,l}^{{\rm b},u}\sim{\rm Exp}\left(\frac{S_{t}^{u}}{P_{u}}\right),$
    so that $\mathbb{E}[T_{t,l}^{{\rm b},u}]=\frac{S_{t}^{u}}{P_{u}}$.
  2. B2
    Communicational Model: We model the per-round communication time $T_{t}^{{\rm c},u}$ as a user-dependent, deterministic value that is constant across rounds:
    $T_{t}^{{\rm c},u}=B_{u}$
  3. B3
    Batch Size: The batch size of user $u$ in round $t$ is dictated by a global scaling parameter $m$, such that:
    $S_{t}^{u}=\Bigg\lfloor mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)\Bigg\rfloor.$
Model Formulation B1 represents the computation time of a layer as an exponential random variable. The exponential model is memoryless and well-suited for modeling compute time variability in deadline-constrained execution. This assumption aligns with prior work in straggler modeling and distributed training analysis [41, 42, 43], where exponential or shifted-exponential distributions are used to enable tractable performance bounds and scheduling policies. Model Formulation B2 models the communication time for each client as a deterministic constant. This is justified by the fact that the transmission time is dominated by stable bandwidth and latency characteristics of the client’s connection. While network fluctuations exist in practice, their variance is typically negligible compared to the overall training timescale [44]. Model Formulation B3 ensures that faster devices do not dominate training by contributing excessively more updates within the same deadline, which could bias the model toward their data and hinder convergence to the global optimum ${\bm{w}}_{\rm opt}$ in (1) [18, 31, 45, 46].
In addition, we define the heterogeneity gap, quantifying the differences in data distribution between users [18]:
$\Gamma\triangleq F\left({\bm{w}}_{\mathrm{opt}}\right)-\frac{1}{U}\sum_{u=1}^{U}\min_{{\bm{w}}}F_{u}({\bm{w}}),$ (6)
where ${\bm{w}}_{\rm opt}$ is defined in (1). A list of notation and symbols used throughout the analysis, including those defined in the preceding sections, is provided in Table I.
TABLE I: Core Notation for ADEL-FL
Symbol Description Defined in
FedAvg
$U,u$ index of the user
$R,t$ index of the global round R1
$L,l$ index of the DNN layer
$T_{\max}$ total training time budget R2
$T_{t}^{{\rm b},u}$ backpropagation time for user $u$ for round $t$
$T_{t}^{{\rm c},u}$ communication time for user $u$ for round $t$
$T_{t}^{u}$ total iteration time for user $u$ for round $t$
$S_{t}^{u}$ batch size of the $u$th client at time $t$
$i_{t}^{u}$ data sample index of the $u$th client at time $t$
$\eta_{t}$ learning rate at global round $t$
$\nabla F_{u}$ stochastic gradient of the $u$th client
${\bm{w}}_{\rm opt}$ global minimizer of the cost function Eq. (1)
${{\bm{w}}}_{t}$ FedAvg global model for round $t$ Eq. (3)
ADEL-FL
$T_{t}^{\rm d}$ adaptive per-round deadline for round $t$
$m$ global batch-scaling parameter B3
$\mathcal{U}_{t}^{l}$ set of users contributing an update for layer $l$ Eq. (4)
$p_{t}^{l}$ probability that $|\mathcal{U}_{t}^{l}|=0$ Eq. (5)
$\tilde{{\bm{w}}}_{t}^{l}$ ADEL-FL global layer $l$ update for round $t$ Eq. (5)
${\bm{w}}^{l}_{u,t}$ ADEL-FL user $u$ layer $l$ update for round $t$
Analysis Parameters
$\rho_{s},\rho_{c}$ smoothness and convexity constants A1
$\sigma_{u}^{2},G$ variance and gradient bounds A2, A3
$P_{u}$ computational capability of user $u$ B1
$B_{u}$ communication time for user $u$ B2
$\Gamma$ heterogeneity gap Eq. (6)
III-B2 Convergence Bound
Under the above assumptions and model formulations, we can now derive the bias correction factor $p_{t}^{l}$ defined in (5):
@@ III-B2 Convergence Bound
When Assumptions A1-A3 and Model Formulations B1-B3 hold, the value of $p_{t}^{l}\triangleq P(|\mathcal{U}_{t}^{l}|=0)$ is bounded by:
$p_{t}^{l}\leq Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U},$ (7)
where $Q(s,x)\triangleq\frac{1}{\Gamma(s)}\int_{x}^{\infty}t^{s-1}e^{-t}dt$ is the regularized upper incomplete gamma function [47].
As expected, $p_{t}^{l}$ is monotonically decreasing with the layer index $l$, as the recursive structure of the backpropagation algorithm implies that first layers are less likely to be reached under tight deadline. Using Lemma 1, we state two lemmas for unbiasedness and bounded variance of each round of our proposed framework:
@@
(Unbiasedness). Under Assumptions A1-A3 and Model Formulations B1-B3, for every FL round of index $t$ and a given set of training data samples $\{i_{t}^{u}\}$, the global model aggregated via ADEL-FL using (5) is an unbiased estimator of the one obtained via vanilla FedAvg [18], namely,
$\mathbb{E}\left[\tilde{{\bm{w}}}_{t+1}\right]={\bm{w}}_{t+1},$ (8)
where ${\bm{w}}_{t+1}$ is given by (3).
@@
(Bounded variance) Consider ADEL-FL with Assumptions A1-A3, Model Formulations B1-B3 and a given set of training data samples $\{i_{t}^{u}\}$. Given the learning rate $\eta_{t}$ set to be non-increasing and satisfying $\eta_{t}\leq 2\eta_{t+1}$, $T_{t+1}^{\rm d}\leq T_{t}^{\rm d}$ and $p_{t}^{1}<0.2$, then, for every FL round $t$, the expected difference between $\tilde{{\bm{w}}}_{t+1}$ and ${\bm{w}}_{t+1}$, which is the variance of $\tilde{{\bm{w}}}_{t+1}$, is bounded by:
$\begin{gathered}\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{t+1}-{\bm{w}}_{t+1}\right\|^{2}\right]\leq\eta_{t}^{2}G^{2}\frac{4U}{(U-1)}\times\\ \sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}{1-5Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}.\end{gathered}$ (9)
Lemmas 2 and 3 are critical for establishing the convergence bound. Lemma 2 ensures that ADEL-FL update remains an unbiased estimator of the FedAvg update, preserving convergence direction. Lemma 3 guarantees that the variance introduced by partial updates remains controlled. These properties mirror the core assumptions required in classical FL convergence analyses, such as those in [18], and are necessary to derive a meaningful upper bound on the expected distance to ${\bm{w}}_{\rm opt}$. We are now ready to state the main theoretical contribution of this work: a novel convergence bound for the ADEL-FL framework that, for the first time, characterizes the fundamental interplay between adaptive per-round deadlines and batch-size scaling. Unlike prior frameworks that treat round timing and local workload as static or decoupled parameters, our result provides a unified analytical objective that captures the trade-off between stochastic gradient noise and deadline-induced truncation variance.
@@
Consider ADEL-FL with Assumptions A1-A3, Model Formulations B1-B3 and a given set of training data samples $\{i_{t}^{u}\}$. Given the learning rate $\eta_{t}$ is set to be non-increasing and holds $\eta_{t}\leq 2\eta_{t+1}$, $\eta_{1}\leq\frac{1}{4\rho_{s}}$, $T_{t+1}^{\rm d}\leq T_{t}^{\rm d}$ and $p_{t}^{1}<0.2$. Then, we have:
$\begin{gathered}\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{R+1}-{\bm{w}}_{\rm opt}\right\|^{2}\right]\leq\prod_{t=1}^{R}\left(1-\eta_{t}\rho_{c}\right)\Delta_{1}\\ +\sum_{t=1}^{R}\eta_{t}^{2}\left(B_{t}+C_{t}\right)\prod_{\tau=t+1}^{R}\left(1-\eta_{\tau}\rho_{c}\right),\end{gathered}$ (10)
where $\Delta_{1}\triangleq\mathbb{E}\left[\left\|\tilde{{\bm{w}}}_{1}-{\bm{w}}_{opt}\right\|^{2}\right]$, and $B_{t}$ and $C_{t}$ are given by:
$B_{t}$ $\triangleq\frac{1}{U^{2}}\sum_{u=1}^{U}\frac{\sigma_{u}^{2}}{mP_{u}\left(\frac{T_{t}^{\rm d}-B_{u}}{T_{t}^{\rm d}}\right)-1}+6\rho_{s}\Gamma$ (11)
$C_{t}$ $\triangleq G^{2}\frac{4U}{U-1}\sum\limits_{l=1}^{L}\frac{1+Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}{1-5Q\left(L+1-l,\frac{T_{t}^{\rm d}}{m}\right)^{U}}.$
The convergence bound is governed by the constants $B_{t}$ and $C_{t}$, which represent the stochastic gradient variance and the deadline-induced truncation variance, respectively. Specifically, $B_{t}$ captures errors originating from data heterogeneity and finite batch sampling, while $C_{t}$ accounts for the information loss when users fail to complete deep network layers within the allotted time. These terms exhibit a fundamental tension mediated by the batch-scaling factor $m$: a larger $m$ reduces $B_{t}$ by improving update quality, but simultaneously exacerbates $C_{t}$ because the increased computational load leads to higher layer-wise truncation under a fixed deadline. This trade-off is analyzed in detail in Section III-D and serves as a principled, tractable objective for the adaptive scheduling framework formulated in Problem 1.
III-C ADEL-FL Algorithm
Having derived the cost function in Theorem 1, we are now ready to present the complete ADEL-FL algorithm, summarized in Algorithm 1. In the beginning of the learning process (Line 2), the server solves the global optimization problem derived from Theorem 1 to jointly determine the round-wise deadlines $\{T_{t}^{\rm d}\}_{t=1}^{R}$ and the global batch-scaling parameter $m$:
@@ III-C ADEL-FL Algorithm
$\min_{\{T_{t}\}_{t=1}^{R},m}$ $\prod_{t=1}^{R}\left(1-\eta_{t}\rho_{c}\right)\Delta_{1}$
$+$ $\sum_{t=1}^{R}\eta_{t}^{2}\left(B_{t}+C_{t}\right)\prod_{\tau=t+1}^{R}\left(1-\eta_{\tau}\rho_{c}\right)$
s.t. $\sum_{t=1}^{R}T_{t}^{\rm d}\leq T_{\max}.$
$T_{t+1}^{\rm d}\leq T_{t}^{\rm d},\quad\forall t\in\{1,\dots,R-1\},$
@@
The server solves Problem 2 using a trust-region method [48], an iterative strategy well-suited for non convex constrained optimization. Trust-region approaches approximate the objective by a local quadratic model and restrict updates to a neighborhood where the model is deemed reliable. The region’s size is adaptively adjusted based on how well the approximation predicts true improvement. Subsequently, the server computes each user’s batch size $S_{t}^{u}$ via Assumption B3 proportional to their computational capability (Line 3). Once the initial configuration is complete, the server transmits the initial global model ${\bm{w}}_{1}$, the round-specific deadlines $T_{t}^{\rm d}$, and the local batch sizes $S_{t}^{u}$ to each participating client.
In each learning round $t$, users perform local training within their time limit, computing gradients up to a layer $d_{t}^{u}$ based on how far they progress before the effective deadline $T_{t}^{\rm d}-B_{u}$ (Line 6), and then transmit those partial updates to the server (Line 7). Following the expiration of the deadline, the server initiates a layer-wise aggregation process to form the new global model update. For each layer $l\in\{1,\dots,L\}$, the server identifies the subset of participating users $\mathcal{U}_{t}^{l}$ who successfully transmitted their gradients for that specific layer before the cutoff (Line 10). The aggregation is performed using the bias-corrected rule in Eq. (5), which explicitly compensates for the partial participation induced by the deadline to ensure the update remains an unbiased estimator of the full gradient (Line 11). This procedure repeats for $R$ rounds, and the final model is returned in Line 14.
@@
1:Input: Number of rounds $R$, training deadline $T_{\max}$, learning rates $\left\{\eta_{t}\right\}_{t=1}^{R}$, computational capabilities $\{P_{u}\}_{u=1}^{U}$, communication times $\{B_{u}\}_{u=1}^{U}$ and initial model ${\bm{w}}_{1}$
2:  Solve Problem 2 to find $\{T_{t}^{\rm d}\}_{t=1}^{R},m$
3:  Set batch sizes $S_{t}^{u}$ for each $u\in\{1,\ldots,U\}$ and $t\in\{1,\ldots,R\}$ via Model Formulation B3
@@
5:    Clients - do in parallel for each $u$, until deadline $T_{t}^{\rm d}-B_{u}$:
@@
8:    Server - do after deadline time $T_{t}^{\rm d}$:
@@
The overall strategy of ADEL-FL refines the convergence guarantees derived earlier by optimally tuning the time-budgeted training configuration. By solving Problem 2, we adaptively allocate deadlines and batch sizes to each round, enabling balanced participation and efficient learning. The resulting algorithm demonstrates improved performance under strict time constraints, as empirically demonstrated in Section IV.
III-D Discussion